You need a(n) 10% acid solution, which you'll get by adding a(n) 15% acid solution to 18 liters of a(n) 7% acid solution. How many liters must be added?
The amount of the 15% solution to be added is?

Answer by mananth(16946)

--- percent ---------------- Amount
Solution I 7 ---------------- 18 liters
solution II 15 ---------------- x liters
Mixture 10 ---------------- 18 + x liters

7 * 18 + 15 x = 10 ( 18 + x )
126 + 15 x = 180 + 10 x
15 x -10 x = -126 + 180
5 x = 54
/ 5
x = 10.8 liters solution II

m.ananth@hotmail.ca